OpenStudy (cloverracer):

a model rocket is launched from a roof into a large field. the path of the rocket can be modeled by the equation y = -0.06x^2 + 9.6x + 5.4 where x is the horizontal distance, in meters, from the starting point on the roof and y is the height, in meters, of the rocket above the ground. How far horizontally from its starting point on the roof and y is the height, in meters, of the rocket above the ground. How far horizontally from its starting point, will the rocket land? Round your answer to the nearest hundredth. 4.30 m 160.56 m 160.23 m 13.94 m

OpenStudy (cloverracer):

please help I have been waiting for an hour :((

OpenStudy (anonymous):

To find where the rocket lands, make y = 0 y = -.0.06x^2 + 9.6x + 5.4 = 0. This is a quadratic equation.Use the quadratic formula to get x. Find D = b^2 - 4ac = 92,16 + 12.96 = 79.25 -> VD = 9.66 x = -b/2a + VD/2a = 9.6/0.12 + 9.67/0.12 = 160. 58

OpenStudy (cloverracer):

thank you!